OpenStudy (anonymous):

Which is the next term in the number pattern: 77, 64, 53, 44, 37

OpenStudy (anonymous):

32 77-64= 13 64-53= 11 53-44= 9 44-37= 7 37-32=5

OpenStudy (anonymous):

Your Welcome